ULTRA-HIGH RESOLUTION SYNCHROTRON-BASED VUV SPECTROSCOPY USING A NEW FOURIER TRANSFORM SPECTROMETER: FIRST APPLICATION TO THE ABSORPTION SPECTRA OF THE RARE GASES.

\underline{MOURAD ROUDJANE}, NELSON DE OLIVEIRA, LAURENT NAHON, Synchrotron SOLEIL, L’orme des Merisiers, St. Aubin Bp 48, 91192 Gif sur Yvette Cedex, France..

A newly-developped Fourier Transform spectrometer (FTS), based upon a wave-front division interferometer, has recently been connected to the VUV beamline DESIRS on the French synchrotron facility SOLEIL. This instrument, unique in the world, using the pseudo-white coherent background continuum of an undulator, allows to perform ultra-high resolution absorption spectroscopy in the 140 - 40 nm spectral range^{a}^{b}, with a theoretical resolving power reaching 1000 000 over the full range. With the intrinsic multiplex capability of FTS over a wide spectral range (typically 5 % of the central wavelength) it should be a unique tool for the study of high lying electronic states of small diluted matter samples, in particular of atmospheric and astrophysical interest. As a first application and performance test of this new spectrometer, we have re-investigated at ultra-high resolving power (≈ 700 000) the VUV absorption spectrum of Ne, Ar, Kr and Xe rare gases involving Rydberg series in the spectral range [54-104] nm, using a windowless room-temperature absorption cell. The comparison of our first results with previous data for these rare gases^{c}^{d}^{e}^{f} led to the identification of several new transitions between the ground state and the five Rydberg ns,ns’ and nd,nd’ series and to a refined determination of the autoionization parameters of the ns’ and nd’ series.

aDe Oliveira, N.;Joyeux, D.; Phalippou, D.; Rodier, J.C.; Nahon, L.; Polack,F.,(to be published).

bDe Oliveira, N.; Joyeux, D.; Phalippou, D.; Rodier, J.C.; Nahon, L.; Polack,F.; Vervloët, M. AIP Conference Proceedings, Vol 879, pp. 447-450 (2007).

cK. Ito, K. Ueda, T. Namioka, K. Yoshino and Y. Morioka, J. Opt. Soc. Am. B5, 10 (1988).

dK. Yoshino. J. Opt. Soc. Am. 65, 9 (1970).

eL. Minnhagen, J. Opt. Soc. Am. 63, 1185 (1973).

fE. B. Saloman, Phys. Chem. Ref. Data., 33. 33 (2004).